### v2

@@ Abstract
Extracting structured computational representations of historical events from narrative text remains computationally expensive when constructed manually. While RDF/OWL reasoners enable graph-based reasoning, they are limited to fragments of first-order logic, preventing deeper temporal and semantic analysis. This paper addresses both challenges by developing automatic historical event extraction models using multiple LLMs (GPT-4, Claude, Llama 3.2) with three enhancement strategies: pure base generation, knowledge graph enhancement, and Retrieval-Augmented Generation (RAG). We conducted comprehensive evaluations using historical texts from Thucydides. Our findings reveal that enhancement strategies optimize different performance dimensions rather than providing universal improvements. For coverage and historical breadth, base generation achieves optimal performance with Claude and GPT-4 extracting comprehensive events. However, for precision, RAG enhancement improves coordinate accuracy and metadata completeness. Model architecture fundamentally determines enhancement sensitivity: larger models demonstrate robust baseline performance with incremental RAG improvements, while Llama 3.2 shows extreme variance from competitive performance to complete failure. We then develop an automated translation pipeline converting extracted RDF representations into Coq proof assistant specifications, enabling higher-order reasoning beyond RDF capabilities including multi-step causal verification, temporal arithmetic with BC dates, and formal proofs about historical causation. The Coq formalization validates that RAG-discovered event types represent legitimate domain-specific semantic structures rather than ontological violations. This work establishes that optimal enhancement strategy selection requires use-case-specific evaluation criteria rather than one-size-fits-all approaches, while demonstrating how automatic ontology construction can be grounded in formal verification for computational humanities applications. [^1] Code and experimental data are available at https://github.com/StergiosCha/RAGGed_events.
@@ 1 Introduction
These findings reveal that enhancement strategies optimize different performance dimensions rather than providing universal improvements. Base generation excels in comprehensive historical coverage and narrative richness, while RAG enhancement trades breadth for technical precision—improving coordinate accuracy and metadata completeness at the cost of event volume and semantic coherence. This trade-off suggests that optimal strategy selection must be driven by specific use-case requirements rather than blanket enhancement application.
To address the fundamental expressivity limitations of RDF/OWL systems revealed through our extraction analysis, we develop an automated translation pipeline that converts extracted knowledge representations into formal specifications for the Coq proof assistant. This translation enables higher-order reasoning capabilities impossible within traditional semantic web frameworks, including verification of multi-step causal relationships, complex temporal arithmetic with historical dates, and formal proofs about historical causation. The Coq formalization also validates a crucial theoretical finding: RAG-discovered event types that appear as ontological violations when evaluated against general-purpose knowledge graphs actually represent legitimate domain-specific semantic structures amenable to rigorous mathematical proof.
Our methodology employs historical texts from Thucydides’ History of the Peloponnesian War as a controlled domain for systematic evaluation. This choice enables assessment of enhancement strategies against well-documented historical events while revealing how model architecture and domain characteristics interact to determine optimal knowledge integration approaches. The integration of automatic extraction with formal verification establishes a foundation for computational humanities applications that combine modern NLP scalability with mathematical rigor. The central contribution of this work is demonstrating that enhancement strategy effectiveness depends fundamentally on the interaction between model capability, domain characteristics, and intended application requirements, challenging the field’s assumption that more comprehensive external knowledge necessarily leads to better performance.
@@ 3 Methodology
In phase 1, we use three enhancement strategies across three LLM architectures using historical text from Thucydides’ History of the Peloponnesian War (first 9 chapters). Our knowledge sources include the DBpedia ontology that involves 768 classes and 3,000+ properties, and the specialized LACRIMALit ontology) for Ancient Greek geographical locations [27]. External knowledge retrieval includes the Wikidata and the DBpedia SPARQL endpoints, as well as the ConceptNet API. The three models we use are GPT-4o (large commercial model), Claude-3.5 Sonnet (alternative large-scale architecture), and Llama 3.2 (3.2B parameter model via local Ollama deployment). All models use deterministic settings (temperature=0) for consistency and use the same chunking parameters. We use three variations on these models, base-generation, generation with knowledge enhancement, and generatoin with knowledge enhancement and RAG.
True Base Generation (Control): Pure LLM inference without external knowledge augmentation. All APIs, knowledge graphs, location enrichment, and RAG functionality are explicitly disabled for baseline measurement. However, the baseline uses the structured prompt shown in figure 1.
@@ 3.2 Enhancement Strategy Comparison
Table 1: Model Performance Comparison
@@ 4 Beyond RDF: Enabling Higher-Order Reasoning
While our event extraction system successfully produces structured RDF knowledge graphs from historical texts, these representations encounter fundamental computational barriers when subjected to sophisticated analytical demands. The intricate temporal dependencies, multi-layered causal networks, and emergent narrative patterns inherent in historical discourse require reasoning architectures that transcend the computational boundaries of conventional semantic web frameworks.
@@ 4.2 Formal Translation Architecture
The translation system comprises four integrated components: (1) an RDF parsing module that extracts event structures, temporal relationships, and participant roles from Turtle serializations; (2) a type inference engine that maps RAG-discovered event classes to formal Coq inductive types, demonstrating that fine-grained classifications like MigrationEvent and WarEvent represent legitimate semantic structures rather than ontological violations; (3) a temporal logic translator that converts chronological relationships into mathematical constraints using dependent types; and (4) a theorem generation system that produces verifiable propositions about historical causation, enabling formal verification of complex historical hypotheses.
@@ 4.4 Translation Process
Stage 1: Ontological Discovery Validation. The translation process begins by analyzing the class hierarchy extracted from RDF triples, distinguishing between standard DBpedia classes and RAG-discovered event types. Our validation algorithm assesses each non-standard class using semantic indicators (presence of domain-specific terms like “War”, “Migration”, “Siege”) and contextual coherence (temporal consistency, specific participants, geographical anchoring). This validation step demonstrates that apparent ontological violations often represent legitimate domain knowledge rather than hallucinations—for instance, ET_MigrationEvent represents historically accurate fine-grained event classification that exceeds DBpedia’s granularity.
@@
Stage 4: Theorem Generation. The final stage generates formal theorems that validate both the semantic coherence of RAG discoveries and enable complex historical reasoning. The rag_discoveries_coherent theorem formally proves that RAG-extracted events with proper temporal, participant, and class structure represent valid domain knowledge, while the another theorem demonstrates how domain-specific classifications enhance causal inference capabilities.
This translation architecture validates our central claim: RAG-discovered event types that appear as ontological violations in RDF evaluation can potentially represent valuable semantic innovations that formal reasoning systems can exploit for sophisticated historical analysis beyond the capabilities of traditional semantic web frameworks.
@@ 4.5 Toward Practical Applications: Foundation for Future Development
5 Conclusions
Our systematic comparison of RAG configurations reveals a fundamental insight: optimal RAG design challenges conventional assumptions by demonstrating that base generation without external enhancement often outperforms both moderate and maximal retrieval strategies. The discovery that pure inferential generation achieves superior overall performance compared to enhanced RAG configurations, particularly for well-documented domains, has significant implications for the field.
These findings challenge the prevailing assumption that more comprehensive retrieval necessarily leads to better performance. Our analysis demonstrates that enhancement strategies optimize different performance dimensions rather than providing universal improvements. Stronger models like GPT-4o and Claude-3.5 demonstrate superior performance in pure base generation, producing comprehensive historical coverage with rich semantic detail, while external enhancement provides complementary but often detrimental effects: improved technical precision at the cost of reduced event coverage and introduction of hallucinations.
The architectural enhancement divide we identified suggests that model capacity fundamentally determines both baseline performance and enhancement sensitivity. Larger, extensively trained models possess robust internal knowledge representations that benefit selectively from external enhancement while potentially sacrificing narrative breadth. Smaller models like Llama 3.2 require external knowledge scaffolding but exhibit catastrophic failure modes when enhancement complexity exceeds their architectural robustness, demonstrating high variance from competitive performance to complete semantic breakdown.
@@ 5 Conclusions
Beyond performance metrics, our work reveals that RAG systems perform ontological discovery rather than mere information retrieval. The fine-grained event types that appear as violations when evaluated against general-purpose ontologies like DBpedia can actually represent legitimate domain-specific semantic structures. Events such as ET_Revolt or ET_MigrationEvent exceed DBpedia’s granularity while maintaining historical accuracy and semantic coherence. This reframes the evaluation paradigm from conformity to existing knowledge graphs toward recognizing domain-specific innovations.
Our translation to Coq demonstrates that these RAG-discovered event knowledge bases can support sophisticated logical reasoning impossible within traditional semantic web frameworks. The ability to express and verify multi-step causal relationships, perform complex temporal arithmetic with BC dates, and conduct formal proofs about historical causation represents a significant advancement beyond RDF/OWL capabilities. The successful integration of RAG-discovered event types into Coq’s rich type system validates their semantic legitimacy through mathematical proof.
The methodology establishes a foundation for computational humanities applications that combine the scalability of modern NLP techniques with the rigor of formal logical systems. While current implementation requires substantial formal logic expertise, the architecture proves that automatic ontology construction can be grounded in mathematical verification, opening promising directions for future tool development.
These findings suggest that RAG system design should prioritize careful evaluation of whether external enhancement is necessary, with configuration choices informed by domain-specific characteristics and model capabilities. For classical historical texts, the LLM’s built-in knowledge proves remarkably comprehensive, suggesting that the value proposition of external enhancement requires careful domain-specific evaluation rather than blanket application.
Future work should explore generalization across domains and historical periods, investigate hybrid approaches combining base generation’s comprehensive coverage with targeted enhancement’s precision, and develop accessible interfaces for formal verification. The integration of ontological discovery validation with automated reasoning offers promising directions for computational linguistics and digital humanities.

### v3

@@ Abstract
Extracting structured representations of historical events from narrative sources still remains challenging when one constructs them manually. While RDF/OWL reasoners support graph-based reasoning, their expressiveness is limited to restricted fragments of first-order logic. We develop automated models for historical event extraction using large language models (GPT-4, Claude, Llama 3.2) with three strategies: direct generation, knowledge-graph augmentation, and retrieval-augmented generation (RAG). Using the 10 first chapters of Thucydides works as a case study, we find that different enhancement strategies optimize different performance dimensions rather than providing across the board universal improvements. Direct generation favors coverage, while RAG improves precision but reduces breadth. Model architecture influences this trade-off: large models show stable baselines with incremental RAG benefits, while Llama 3.2 exhibits extreme variance from competitive to catastrophic performance. To address RDF’s expressivity limitations, we develop a translation pipeline converting RDF outputs to Coq proof assistant specifications, enabling temporal arithmetic with BCE dates, multi-step causal inference, and formal validation of domain-specific event types. This demonstrates that optimal enhancement strategies depend on specific application requirements, while establishing foundations for computational humanities combining NLP scalability with formal verification. [^1] Code and experimental data are available here: https://drive.google.com/drive/folders/1ZkZc3CnNGAdXLUzKly4-FGXiVqk9BpaW?usp=sharing .
@@ 1 Introduction
These findings reveal that enhancement strategies optimize different performance dimensions rather than providing universal improvements. Base generation excels in comprehensive historical coverage and narrative richness, while RAG enhancement trades breadth for technical precision-improving coordinate accuracy and metadata completeness at the cost of event volume and semantic coherence. This trade-off suggests that optimal strategy selection must be driven by specific use-case requirements rather than blanket enhancement application.
To address the fundamental expressivity limitations of RDF/OWL systems revealed through our extraction analysis, we develop an automated translation pipeline that converts extracted knowledge representations into formal specifications for the Coq proof assistant. This translation enables higher-order reasoning capabilities impossible within traditional semantic web frameworks, including verification of multi-step causal relationships, complex temporal arithmetic with historical dates, and formal proofs about historical causation. The Coq formalization also validates a crucial theoretical finding: RAG-discovered event types that are missing when evaluated against general-purpose knowledge graphs actually represent legitimate domain-specific semantic structures amenable to rigorous mathematical proof.
Our methodology employs historical texts from Thucydides’ History of the Peloponnesian War as a controlled domain for systematic evaluation. More specifically, we use the ten first chapters of Thucydides in the English translation by Richard Crawley This choice enables assessment of enhancement strategies against well-documented historical events while revealing how model architecture and domain characteristics interact to determine optimal knowledge integration approaches. The integration of automatic extraction with formal verification establishes a foundation for computational humanities applications that combine modern NLP scalability with mathematical rigor. The central contribution of this work is demonstrating that enhancement strategy effectiveness depends fundamentally on the interaction between model capability, domain characteristics, and intended application requirements, challenging the field’s assumption that more comprehensive external knowledge necessarily leads to better performance.
@@ 2 Related work
Rather than contradicting the general insight that accurate, targeted knowledge improves LLM performance, our experiments refine it: we show that injection of excessive or loosely filtered context can degrade extraction quality, especially in already knowledge-rich models. We call this the inverse calibration principle.
@@ 3 Methodology
In phase 1, we use three enhancement strategies across three LLM architectures using historical text from Thucydides’ History of the Peloponnesian War (first 9 chapters). Our knowledge sources include the DBpedia ontology that involves 768 classes and 3,000+ properties, and the specialized LACRIMALit ontology) for Ancient Greek geographical locations [27]. External knowledge retrieval includes the Wikidata and the DBpedia SPARQL endpoints, as well as the ConceptNet API. The three models we use are GPT-4o (large commercial model), Claude-3.5 Sonnet (alternative large-scale architecture), and Llama 3.2 (3.2B parameter model via local Ollama deployment). All models use deterministic settings (temperature=0) for consistency and use the same chunking parameters. We use three variations on these models, base-generation, generation with knowledge enhancement, and generation with knowledge enhancement and RAG.
True Base Generation (Control): Pure LLM inference without external knowledge augmentation. For this approach, the input text is segmented into chunks, and the model processes these chunks one by one, relying solely on its internal knowledge to complete the extraction. All APIs, knowledge graphs, location enrichment, and RAG functionality are explicitly disabled for baseline measurement. However, the baseline uses the structured prompt shown in figure 1.
@@
The placeholders {entities} and {locations} are automatically populated by a preprocessing pipeline: named entities are extracted via regex/NER heuristics (limit 25 per chunk) and candidate locations are identified through a LocationExtractor that consults regex patterns, the LACRIMALit ontology, and then Wikidata/DBpedia. RAG chunks are obtained via FAISS with MiniLM embeddings; we used chunk size 1,500–2,000 tokens with 200 character overlap. Evaluation combined manual expert assessment of correctness with quantitative counts of extracted events, coordinate precision (Geo distance from known sites), and RDF formatting completeness.
@@ 3.2 Enhancement Strategy Comparison
Table 1: Model Performance Comparison. Here “No RAG + KG” refers to enrichment from a single source (primary triple retrieval), while “RAG + Multi KG” combines context from all three sources (Wikidata, DBpedia, ConceptNet) with vector similarity retrieval.
@@ 4 Beyond RDF: Enabling Higher-Order Reasoning
While our event extraction system successfully produces structured RDF knowledge graphs from historical texts, these representations encounter fundamental computational barriers when subjected to more fine-grained demands. Our RAG evaluation demonstrates that no matter the enhancement strategy, be it base generation, knowledge augmentation, or full RAG, the RDF knowledge graphs we get identical limitations when processed by conventional reasoning systems. More fine-grained temporal dependencies, multi-layered causal networks, and emergent narrative patterns common in historical discourse require reasoning architectures that can extend beyond systems, albeit extremely useful ones, that are limited to fragments of first-order logic.
@@ 4.1 Computational Limitations of RDF/OWL Systems
It is true that extensions such as SWRL or DL-safe Horn rules can cover additional entailed facts beyond pure RDFS. However, these rule regimes remain essentially first-order, lacking facilities for dependent types, integer date arithmetic over BCE conventions, or higher-order causal composition, needed for careful and fine-grained reasoning over events. Our use of Coq thus complements rather than replaces SWRL-style reasoning: SWRL is well-suited for monotonic rule expansions, while Coq enables formal arithmetic and multi-step causal proofs that exceed those capabilities.
@@ 4.2 Formal Translation Architecture
The translation system comprises four integrated components: (1) an RDF parsing module that extracts event structures, temporal relationships, and participant roles from Turtle serializations; (2) a type inference engine that maps RAG-discovered event classes to formal Coq inductive types, demonstrating that fine-grained classifications like MigrationEvent and WarEvent represent legitimate semantic structures and extensions of a pre-existing ontology; (3) a temporal logic translator that converts chronological relationships into mathematical constraints using dependent types; and (4) a theorem generation system that produces verifiable propositions about historical causation, enabling formal verification of complex historical hypotheses.
@@ 4.4 Translation Process
Stage 1: Ontological Discovery Validation. The translation process begins by analyzing the class hierarchy extracted from RDF triples, distinguishing between standard DBpedia classes and RAG-discovered event types. Our validation algorithm assesses each non-standard class using semantic indicators (presence of domain-specific terms like “War”, “Migration”, “Siege”) and contextual coherence (temporal consistency, specific participants, geographical anchoring). This validation step demonstrates that apparent ontological gaps represent legitimate domain knowledge rather than hallucinations—for instance, ET_MigrationEvent represents historically accurate fine-grained event classification that exceeds DBpedia’s granularity.
@@
Stage 4: Theorem Generation. The final stage generates formal theorems that validate both the semantic coherence of RAG discoveries and enable complex historical reasoning. The rag_discoveries_coherent theorem formally proves that RAG-extracted events with proper temporal, participant, and class structure represent valid domain knowledge, while another theorem demonstrates how domain-specific classifications enhance causal inference capabilities.
This translation architecture validates our central claim: RAG-discovered event types that do not appear in the event database but can be used to extend the ontology to represent valuable semantic innovations that formal reasoning systems can exploit for sophisticated historical analysis beyond the capabilities of traditional semantic web frameworks.
@@ 4.5 Toward Practical Applications: Foundation for Future Development
5 Conclusions, Limitations and Future Work
Our systematic evaluation of RAG-enhanced historical event extraction reveals that enhancement strategy effectiveness depends fundamentally on model architecture and domain characteristics rather than following universal patterns. Stronger models like GPT-4o and Claude-3.5 achieve optimal performance through base generation (extracting 36 and 39 events respectively), while external enhancement often proves detrimental—introducing hallucinations and reducing coverage. Conversely, weaker models like Llama 3.2 benefit from controlled knowledge augmentation but exhibit catastrophic failure under complex RAG configurations, demonstrating an inverse calibration principle where model capability inversely correlates with enhancement benefit.
These findings challenge the assumption that comprehensive retrieval necessarily improves performance. Enhancement strategies optimize different dimensions: base generation excels in coverage and narrative richness, while RAG trades breadth for technical precision. This suggests that optimal strategy selection must be driven by specific application requirements rather than blanket enhancement approaches.
@@ 5 Conclusions, Limitations and Future Work
Beyond extraction optimization, our work reveals that RAG systems discover domain-specific ontological structures that exceed general-purpose knowledge graphs. Fine-grained event types like ET_MigrationEvent and ET_WarEvent, which appear as gaps when evaluated against DBpedia, represent legitimate semantic innovations that capture historically accurate distinctions essential for domain analysis.
To address the expressivity limitations of RDF/OWL systems revealed through our extraction analysis, we developed an automated translation pipeline converting extracted representations into Coq proof assistant specifications. This enables higher-order reasoning impossible within traditional semantic web frameworks: multi-step causal verification, temporal arithmetic with BCE dates, and formal validation of RAG-discovered event types. The successful integration of these domain-specific classifications into Coq’s type system validates their semantic legitimacy through mathematical proof.
A main limitation of this current work is that our evaluation focuses on classical Greek historical texts, providing controlled conditions but limiting generalizability. Future work should systematically evaluate these patterns across diverse historical periods, languages, and narrative styles to establish broader applicability of the inverse calibration principle. Additionally, while our Coq translation demonstrates formal reasoning feasibility, practical deployment requires accessible interfaces that reduce the formal logic expertise barrier for working historians.
This work establishes a foundation for computational humanities that combines NLP scalability with formal verification rigor. For knowledge-rich domains like classical history, the internal knowledge of LLMs is shown to be remarkably comprehensive, suggesting that external enhancement requires careful domain-specific evaluation rather than universal application. Lastly, the integration of automatic extraction with formal reasoning opens promising directions for tools that can support both large-scale processing and rigorous analytical workflows demanded by historical scholarship.